# Astrobites Collaboration Improving Undergraduate Astronomy Students’ Skills with Research Literature via Accessible Summaries: A Case Study with Astrobites-based Lesson Plans

### v2

@@ III.1 Observed Changes in Student Attitudes
The analysis for the conceptual categories of questions, as summarized in Table 1 and provided in full in Section B, is shown in Table 3. Every category showed significant change from pre- to post-lesson. The effect sizes were small as measured by Cohen’s d, but BF10 provided evidence in favor of change for all groups. Notably, there is extremely strong evidence for change in the groups perceived ability with jargon and main takeaways. There is strong or very strong evidence for perceived ability with conceptual understanding/intuition, perception of science communication ability, and perceived ability within astronomy. Evidence for changes in feelings of belonging within astronomy were only anecdotal. This lack of change in feelings of belonging may be related to Question 26, as described above.
Group | p-value | CI95% | Cohen’s d | BF10 | BF10 | power |
interpretation | ||||||
Jargon | [0.12 0.34] | 0.155 | 118 | extreme | 0.87874 | |
Main Takeaways | [0.16 0.37] | 0.218 | extreme | 0.99799 | ||
Conceptual | [0.09 0.27] | 0.142 | 64.1 | very strong | 0.92311 | |
Communication | [0.07 0.25] | 0.120 | 14.0 | strong | 0.75953 | |
Ability | [0.1 0.33] | 0.163 | 37.3 | very strong | 0.91009 | |
Belonging | [0.04 0.25] | 0.102 | 2.40 | anecdotal | 0.61747 |
Question | CI95% | p | Cohen’s d | d interp. | BF10 | BF10 interp. | Text |
() | for H1 | ||||||
1 | [0.25 0.75] | 1.27 | 0.456 | small | 158 | extreme | I am able to define technical |
terminology that appears in | |||||||
scientific literature. | |||||||
4 | [0.09 0.55] | 0.259 | small | 3.83 | moderate | I am able to understand | |
the technical details | |||||||
of a scientific study. | |||||||
6 | [0.3 0.79] | 0.428 | small | 698 | extreme | After reading a paper, I can | |
summarize it well to a classmate | |||||||
in a science course. | |||||||
24 | [0.22 0.71] | 0.324 | small | 62.1 | very strong | I can explain astronomy concepts | |
to my science instructors. | |||||||
28 | [0.18 0.72] | 0.347 | small | 17.2 | strong | I am able to use scientific literature | |
to guide my studies in astronomy. |
Results from the qualitative coding are summarized in Figures 3 and 4. In response to the question “How do you feel about reading research papers?”, responses citing intimidation lowered by 13.6%, unpreparedness by 6.8%, and enjoyment/fun increased by 6.8%. Additionally, 6.8% more responses described the importance of reading research papers. In response to “What comes to your mind when you think of research?”, all categories show small or no change; students overwhelmingly responded with the topics and skills needed (e.g., coding, observing/telescopes, math, etc.) more than their feelings about it. In response to “Do you think you are capable of being an astronomer?”, responses coded as confident were reduced by 6.5%, but responses coded as insecure were also reduced by 11%. Other codes for this question only show small change, if any. In response to “Do you feel like you belong? Why or why not?”, yes responses slightly increased while no slightly decreased, and maybe responses increased by 4%. Responses citing motivation, experience, and community decreased by 4-5%, while self-doubt also decreased by 2%. All codes for this question show small change, and such minute differences should be interpreted with caution.
Two questions were available on the post-lesson survey only. For the question “What is your main takeaway?”, 44% of responses described an increased understanding of the field, 25% described an understanding of accessibility (e.g., resources, science communication), 23% claimed increased confidence, 15% described furthering their knowledge of astronomy, and a small number reported increased motivation or decreased confidence. In response to “How have your feelings changed?”, 50% of responses described no change, 27% said their feelings were “improved” in one way or another, 12% cited further enjoyment (e.g., appreciation for science communication, more inspired to do astronomy, more excited to read papers), and 11% described increased meta-knowledge of the field (e.g., how to find resources, appreciation for the work involved, career options). Our consistency analysis, described in detail in Appendix F, shows that open-ended responses tend to track Likert scores for each question; however, the Spearman correlation coefficient is smaller in the post-survey than in the pre-survey. When looking at the changes in individual students, students with initially low scores generally improved, students with initially medium scores generally showed no change or a slight benefit, and students with initially high scores generally stayed the same, but a small number decreased, as summarized in the Sankey diagram in Figure 5.
III.2 Direct Feedback on Lesson Plans
Both students and instructors were asked to provide direct feedback on the lesson plans, including their enjoyment, what they found most useful, and what they found most confusing. Below is a summary of their responses on the lesson plans.
Students had an overwhelmingly positive sentiment towards the lesson plans when asked, “Did you enjoy this activity with Astrobites?”. 103/119 (86.6%) said they enjoyed the lessons, with 12/119 (10%) indifferent and 4/119 not enjoying (3.4%). (Note: for direct feedback, we included responses from students who did not take the pre-lesson survey, making the sample size 119 instead of the 82 for analysis of pre-post pairs.) The four who did not enjoy the activity cited an instructor’s addition to the lesson, personal preference to not read long articles, the lesson plans not being challenging enough, or wanting more freedom to choose which paper they read (versus one assigned by their instructor).
Students found the lessons useful because of the real-world applications of the assignment, increased engagement and discussion in class facilitated by the activities, the opportunity to practice science writing and communication, and the opportunity to get a better sense of career options and role models in the field, as expressed in the below quotes from student responses. The most common response regarding usefulness, though, was the exposure to research literature and practice reading papers provided by the lesson plans; five students explicitly cited Astrobites itself as a useful resource they’re glad to have learned about. Finally, one student encapsulated their main takeaways as follows, going into detail on how this lesson plan benefited them:
“The main takeaway from this activity if that when you take things step by step, they aren’t as daunting as they seem. Learning about the scientist, then a major concept they are involved with, and lastly a paper they partook in was surprisingly easy to understand. I feel accomplished for my understanding of all this new material in less then two weeks and I feel like this information will stick with me for the long run.”
When asked what was most confusing about the lesson plans, students overwhelmingly cited plots and jargon from the papers–exactly what we’d expect for students just learning how to engage with scientific literature, as quoted below. Students did not provide any actionable feedback on the lesson plan itself.
“The scholarly article was the most difficult and confusing, but I am really glad I overcame that and learned what all of that research had offered me. Amazing.”
@@ III.2 Direct Feedback on Lesson Plans
All instructors said the lesson plan was well-received in their class, and mentioned the following benefits of the lesson plans (with illustrative quotes below): helping to get students reading papers, exposing them to what kind of research exists in the field, increased engagement in class, increases in student confidence, awareness of Astrobites as a resource, and a change of pace from a normal routine (e.g., problem sets).
@@
Instructors nearly unanimously stated the most useful aspect of the lesson plan was simply that it exists; structured/pre-made lesson plans connecting to current research that can be easily slotted into a course are an extremely beneficial resource for already over-committed educators. The resource of Astrobites in particular, as described by one educator, enables connections to literature that may not be possible otherwise for an introductory course:
“With an Astrobites post, it’s been since simplified enough that we can cover it in one class, but they still get a lot of the same benefits they get from working directly with the article.”
@@
Instructors additionally provided helpful feedback and suggestions for improvement, which we will take under consideration next time we revise the lesson plan resources. These improvements include: incorporating “classic” (e.g., foundational or historical) papers for their pedagogical value and fundamental roles in the field, adding a structured option to do Lesson Plan 1 (readings) as a recurring activity, and adding a reflection component for the presentations (e.g., “What made you choose this scientist?”). One instructor also requested a forum for sharing Astrobites lesson plans and resources between instructors, which we hope to implement on our website in the near future. Additional illustrative quotes from responses are available in Appendix D.
IV Discussion
Both students and educators expressed positive feelings towards the Astrobites lesson plans in their direct feedback, and our data show positive change in all major categories probed by the quantitative survey questions (see Tables 1), as summarized in Section III.1. The strongest effect sizes were measured for ability with jargon and main takeaways. The majority of students participating in the study experienced Lesson Plan 1 (Reading Assignments), where the learning goals are expressly focused on reading comprehension and jargon. When broken down by lesson plan, the data clearly show that the group experiencing Lesson Plan 1–which had the most participants, and directly focused on reading comprehension skills–was driving the extreme evidence for improvement with main takeaways. Additionally, qualitative responses such as “I generally find it interesting, especially now that I have a base of knowledge to understand the key concepts” and “usually is a daunting task but Astrobites makes it manageable” highlight the usefulness of these lesson plans and the accessible summaries for students’ literature-reading skills.
@@ IV Discussion
The amount of students whose scores began and remained high (as shown in Figure 5) highlights a limitation in our analysis: students whose initial responses lay to one extreme or another (the “floor and ceiling effects” [49]), i.e., students whose initial averages in a category were close to 1 or 7. A student with an initial average of 2 could not exhibit a decrease of more than after experiencing the lesson plans, while a student with an initial average of 6 could not exhibit an increase of more than . It is also possible that the use of Astrobites may have confused initially high-scoring students, leading to the small number of decreases in scores, suggesting that the existing Astrobites lesson plans may be good for introducing students to a topic but less effective for delving into it further. Alternatively, it is important to note we only completed the treatment once, and it make take repeated exposures (i.e. more practice with the lesson plans) for benefits to be realized.
@@
Qualitative responses from students showed a more mixed and nuanced view of their development from before to after experiencing the lesson plans. Feelings of belonging were largely improved in the qualitative responses, although their reasons for their responses varied unpredictably. In response to questions about feelings while reading research, students less often mentioned feelings of intimidation and unpreparedness after the lesson, while mentioning more often that papers feel important/worthwhile and they are confident in their abilities. However, they also more often cited papers as requiring effort, and less often as enjoyable after the lesson; we interpret this as students having a more realistic understanding of the task at hand, for better or worse, while also being better equipped to handle it. Similarly, students reported fewer feelings of insecurity when asked about their ability to be a scientist, but also fewer feelings of confidence. Decreases in confidence may also be attributed to the fact that students’ “measuring stick” for themselves has changed since the beginning of the course, as they have gained knowledge and practice [50, 51].
@@
For a majority of students, as shown in Figure 4, the main takeaway from the lesson was a greater understanding of the field of astronomy, e.g. how research works, what scientists do, how research publications come to be. Nearly all students’ takeaways were positive and in line with the learning objectives from the lesson plans, including increased content knowledge, increased confidence with course material, and a greater awareness of resources and the need for accessibility in science. Although 50% of students claim their feelings about astronomy have not changed as a result of the activity, those who do claim to have changed are remarkably all positive in some way.
@@
Students were generally consistent in their qualitative and quantitative responses for the pre-survey, but this is not the case in the post-survey. We argue that the greater nuance available in written responses meant that students were more likely to express a mix of positive and negative feelings — with our qualitative scoring model for consistency analysis, this could translate to more neutral or negative scores, even if their overall feelings (as reflected through a collection of Likert scale questions) showed positive change.
Additionally, students and instructors were largely consistent in their feedback – instructors reported feeling that their students had gained understanding and perspective from the lessons, and most students reported either improved or reinforced senses of belonging and understanding in astronomy. Feedback also supported that our lesson plans were addressing their intended learning goals, which vary slightly across the four available activities; from post-implementation interviews with the instructors, those who used Lesson Plan 1 (reading assignments) commented more on the usefulness of exposing students to literature, while those who used Lesson Plan 4 (presentations from Astrobites interviews) commented about belonging, diverse role models, and career options.
V Conclusions and Recommendations

### v3

@@ III Quantitative Analysis & Results
The analysis for the conceptual categories of questions is shown in Table 2. Every category except ‘Belonging’ showed significant change (interpreted from the calculated p-values) from pre-lesson to post-lesson; BF10 provided evidence in favor of change for all groups except ‘Communication’ and ‘Belonging’. Note that these two categories have low power and thus true effects may not be detected. Notably, there is extremely strong evidence for change in the group’s perceived ability with main takeaways. There is strong evidence in change in perceived ability with jargon and moderate evidence for change in perceived ability with conceptual understanding/intuition. Evidence for changes in perception of ability within astronomy was only anecdotal. There was a preference for no change in perception of science communication ability and feelings of belonging within astronomy.
Students improved to some degree across all categories, with the highest degree of change in student perception of ability to extract main takeaways; it is worth reiterating that these are student perceptions of their ability with primary research articles, as affected by their engagement with accessible research summaries.
IV Qualitative Analysis & Results
We explored student responses to open-ended survey questions regarding perception of ability with research & literature (Questions 36, 38, 41, and 44 in Appendix B) through thematic analysis [51]. Specifically, we use the coding reliability approach for thematic analysis, wherein we aim for consistency between coders. These four questions were chosen for their relevance to the learning outcomes of the reading comprehension assignment template.
A set of thematic categories (referred to herein as codes, summarized in Table 3) were determined inductively for each qualitative question, based on the content of student responses. All responses were read before creating codes, and then responses were re-read several times to refine the initial set of themes/codes. After collective creation of these codes by the study team, seven raters (authors BLL, NKC, RRL, SLW, SG, MD and collaborator AM) independently assigned codes to each response. It is worth noting that three of the raters (NKC, SLW, AM) were not involved in creating the codebook; this independent group serves as a simple verification for the codes used. We only included and coded responses from the same sample as the quantitative analysis ().
| Survey Question | Codes |
| Q36: How do you feel about reading scientific research papers? | Intimidation (intimidated, daunting, overwhelmed, confused, not comfortable); Require effort (time-consuming, tedious, read multiple times); Worthwhile (rewarding, useful, necessary); Unprepared (lacking knowledge, not taken enough classes, “not ready yet”); Improved (e.g., “better now”); Confidence (comfortable, good, “like I can do it”); Enjoyment(interesting, fun, enjoy) |
| Q38: Do you think you are capable of being an astronomer (or related scientist)? Why or why not? | Yes; No; Maybe (depends, yes and no); Confident (good at it, previous or current experience, have what it takes, success in classes); Insecure/uncertain (no, maybe, unsure, grades, imposter syndrome, doubt); Passion (life long interest, plans for graduate school) |
| Q41: What is your main takeaway from this activity? | Astronomy knowledge (topics, interpreting plots, data); Understanding of field (careers, breadth of topics, better, how papers are written, how science works and results are communicated, programming, statistics, visuals); Confidence (boost, more capable, more confident, less daunting, more daunting, decreased); Accessibility (more accessible, digestible, resources. importance of good teaching and/or accessible literature); Decreased confidence (more daunting, decreased) |
| Q44: How have your feelings about astronomy research changed as a result of this activity, if at all? | Improved (more capable, less daunting, more confident); Enjoyment (appreciation for sci. comm., more inspired to do astro., more excited to read papers); Meta-knowledge of field (awareness of resources, knowledge of career options, accessibility of field); No change |
Due to the inherently subjective nature of coding, we assessed inter-rater agreement using the full set of codes applied by the seven independent coders. While there are a range of statistics used to determine inter-rater agreement and reliability, many of the more commonly used ones (i.e. kappa indices, intraclass correlation coefficient) do not apply to our data wherein we have more than two raters and the ability to apply multiple codes to each response [52]. We instead elected to use the proportion agreement, which we calculate by taking the number of codes for each response that all raters agreed on (either all raters applying the code, or all raters not applying the code) and dividing by the total number of codes it was possible to apply. Thus if there were six codes for a given question, the proportion agreement would be equal to where is the number of codes all seven raters selected for the response and is the number of codes none of the raters selected. This statistic does not take into account the possibility of chance agreement, making it slightly less robust, but given the large sample size and number of coders, as well as the lack of mutual-exclusivity among the codes, we determined this was the best way to proceed. There is a precedent of using proportion agreements in these types of cases [53, 54, 8]. The resulting percentage agreement for each question is summarized in Table 4. In general, while there is no consistently-used cutoff for an acceptable level of agreement, cutoffs generally fall between 0.7 and 0.8, and here we consider a proportion agreement to be acceptable [54, 55, 56, 57]. We calculated these proportion agreements for each question, not per code, due to the nature of our codebook–that is, each question had its own unique codes, and multiple codes were able to be applied to a single response.
| Question Number | Proportion Agreement |
| 36 | 0.778 |
| 38 | 0.783 |
| 41 | 0.717 |
| 44 | 0.815 |
We found that all of the questions used in this analysis fall within the limit of , although Q41 is near the lower end. Upon a brief examination of causes of inconsistency, we found that for responses that did not have complete agreement, agreement levels were still generally high (i.e. 6/7 coders applying a given code). There were no instances where the same response yielded contradicting codes (i.e. both ‘Yes’ and ‘No’ for Q38, or both ‘Improved’ and ‘No change’ for Q44). More often, differences in code applications arose in unpacking the reasoning of participant’s answers. The codes ‘Understanding of field’ in Q41 and ‘Confidence’ in Q36 saw the largest amount of disagreement in applications between coders. These were codes that were usually applied in tandem with other codes, as responses coded ‘Understanding of Field’ were also often coded with ‘Accessibility’, and those coded with ‘Confidence’ also had many coders selecting ‘Passion’. This suggests that individual coders may be drawing some connections between codes not explicitly designated in the codebook. Despite these few instances of lower agreement, the generally strong proportion agreement values allow us to proceed with analyzing these code applications.
It is worth noting that some students left questions blank, or provided responses that were illegible or unable to be matched to a code. Qualitative response codes were tallied, and the number of responses in each main code category (e.g., not including sub-codes) was compared between pre-lesson and post-lesson surveys. For our analysis, codes applied by a majority of raters () were included as the final set of identified codes for each response, essentially determining the interpretation of the response by majority. Codes applied by raters were not included.
IV.1 Student Responses
The raw counts and total differences in number of responses assigned a given code are shared in Tables 5, 6, and 7. In this section, we summarize the notable findings for each qualitative question.
| Q36: How do you feel about reading scientific research papers? | ||||||
| Intimidation | Require Effort | Worthwhile | Unprepared | Confidence | Enjoyment | |
| Pre | 11 | 11 | 7 | 9 | 10 | 17 |
| Post | 14 | 8 | 8 | 2 | 19 | 17 |
| Difference | -7 | -3 | 1 | -7 | 9 | 0 |
| Q38: Do you think you are capable of being an astronomer (or related scientist)? Why or why not? | ||||||
| Yes | No | Maybe | Confident | Passion | Insecure/Uncertain | |
| Pre | 42 | 1 | 8 | 29 | 18 | 12 |
| Post | 45 | 2 | 4 | 22 | 17 | 7 |
| Difference | 3 | 1 | -4 | -7 | -1 | -5 |
| Q41: What is your main takeaway from this activity? | ||||
| Knowledge | Field Understanding | Confidence | Accessibility | Decreased confidence |
| 9 | 14 | 7 | 13 | 3 |
| Q44: How have your feelings about astronomy research changed as a result of this activity, if at all? | ||||
| Improved | Enjoyment | Meta-knowledge | No Change | |
| 11 | 7 | 5 | 25 | |
To analyze the responses to Question 36 (“How do you feel about reading research papers?”) and Question 38 (“Do you think you are capable of being an astronomer?”), we matched students’ pre-survey and post-survey to see how their responses changed over time. It is worth noting that students were writing a new response without access to their pre-intervention survey for the post-intervention survey, so they may have shared different information than on the initial survey; this may account for some of the responses that contained a certain code in the pre-intervention survey but not in the post-intervention survey.
For Q36, 20 of the pre-survey and 16 of the post-survey responses indicated feeling intimidated by or unprepared for scientific papers. There were eight students whose comments suggested feeling intimidated in the pre-survey, but did not in post, although four students were not coded for being intimidated for pre-, but were for post-. It is worth noting that for three out of those four, their post-survey responses also indicated something positive about the experience, as one said “They’re a little daunting, but I usually enjoy it.” There were also eight students who felt unprepared in the pre-survey, but did not mention that concern in the post, and one explicitly stated “Reading astronomy papers has gotten easier with time.” However, four of those eight did still indicate that they were intimidated in their post-survey responses, suggesting they weren’t yet fully comfortable with reading papers even if there was a slight improvement. Many students also stated that they enjoyed reading papers, with thirteen students maintaining that position in both pre- and post- responses, five saying it only in post, and five saying it only in pre. As touched on above, several of these responses came with qualifiers such as “daunting”, “I don’t always understand”, and “sometimes the wording…is hard to follow”. Finally, 10 of the 52 respondents expressed a level of confidence in the post-survey that they did not in the pre-survey, and overall 19 respondents felt at least somewhat confident about reading scientific papers at the end of the intervention. A summary of changes in individual students’ reported feelings towards reading research papers is presented in Figure 2.
For the other paired qualitative question, “Do you think you are capable of being an astronomer?”, a majority of students responded with yes on both the pre-lesson and post-lesson surveys (42 pre, 45 post). Of students whose answers changed from pre-lesson to post-lesson, 6 reported positive change (from “no” or “maybe” to “yes”), 2 reported a slightly negative change (from “yes” to “maybe”), and 1 reported a negative change (from “yes” to “no”). The student whose answer changed to no cited feelings of overwhelm in their post-lesson response: “I don’t think I have the skills necessary and feel quite overwhelmed with the way things work in astrophysics.” Response counts for Question 38 are provided in Table 6, and changes in student responses for Question 38 are illustrated in Figure 3.
Two questions were available on the post-lesson survey only. For Question 41 “What is your main takeaway?”, we categorized student responses into five categories: astronomical knowledge gains (9 students), improvements to students’ understandings of the field/culture of astronomy (14 students), improved confidence (7 students), decreased confidence (3 students), and a greater awareness of the possibilities/need for accessibility in the field (13 students). For the respondents who shared a decreased confidence, the largest cause seemed to be difficulty in reading papers due to the amount of jargon and technical language. As one student put it “Astronomy papers are filled with a lot of information only meant for astronomers.” For those students who had increased confidence, some things they identified that helped them were having a general idea of the scientific background and taking their time with reading. Of the 13 respondents coded for ‘Accessibility’, a common trend in the responses was how the key takeaways from research can be presented more simply than what is seen in papers. As one student explained, “Scientific findings can be phrased in relatively simple ways, making the findings of research more accessible to people who don’t already know a lot about the material.”
For Question 44 “How have your feelings about astronomy research changed?”, eleven students indicated that their feelings improved, 25 indicated that there was no change, and no students reported more negative feelings about astronomy research after the activity. We additionally coded responses for mentions of enjoyment and meta-knowledge of the field, with seven respondents mentioning enjoyment and five gaining more meta-information. Among the meta-knowledge responses, one student mentioned they now knew how to find more resources, while the others focused around more general career advancement, as one respondent shared “I have a better understanding about my future in the field.” Response counts for Questions 41 and 44 are also summarized in Table 7.
IV.2 Instructor Feedback
@@
Due to the small number of instructors, we consider their feedback on all assignment templates implemented, including but not limited to the reading assignment. All instructors said the assignment template they implemented was well-received in their class, and mentioned the following benefits of the assignments: helping to get students reading papers, exposing them to what kind of research exists in the field, increased engagement in class, increases in student confidence, awareness of Astrobites as a resource, and a change of pace from a normal routine (e.g., problem sets).
@@
Instructors nearly unanimously stated the most useful aspect of the assignment template was simply that it exists; structured/pre-made resources connecting to current research that can be easily slotted into a course are an extremely beneficial resource for already over-committed educators.
@@
Instructors additionally provided helpful feedback and suggestions for improvement, which we will take under consideration next time we revise the assignment template resources. These improvements include: incorporating “classic” (e.g., foundational or historical) papers for their pedagogical value and fundamental roles in the field and adding a structured option to do the reading assignment as a recurring activity.Additional illustrative quotes from educator qualitative responses are available in Appendix D.
V Consistency Analysis
When performing a mixed methods study such as this, it is good practice to compare results obtained through multiple questions covering the same themes. One way to conduct such a comparison involves assigning numerical scores to qualitative questions [58, 59].
Here, we identified three qualitative questions (Questions 36, 38, and 39) for which there are Likert scale questions that address similar themes. We determined which qualitative response codes denoted positive, neutral, and negative feelings toward the three questions, and assigned each student a qualitative “score” for that question based on the prevalence of each type of code in their response. The codes considered positive, neutral, and negative for each of the three qualitative questions used in this comparison are listed in Table 8.
| Open-Ended Question | Codes considered | Related Likert questions |
| 36 (Papers) | Intimidation, Unprepared, Effort, | 1, 2, 3, 4, 5, 7, 8, 9, 11, 28 |
| Confidence, Worthwhile, Enjoyment | ||
| 38 (Capable) | Yes, No, Maybe | 14, 18, 25, 26, 27, 28, 29, 30, 31, 32 |
| 39 (Belonging) | Yes, No, Maybe | 32, 33, 34, 35 |
For each student response to each of these three questions, we took the average of related Likert scale questions to find their “Likert score” (quantitative). To assign a student’s “written score” (qualitative) on a topic, we assigned to every positive code applied to their response to that question, to every negative code, and to every neutral code, then applied the average of these scores for that response. We only considered codes applied by a majority of raters, and we considered the number of raters who applied a given code. For example, if four raters applied a positive code to a response but six applied a negative code, the response received a score of . (Note that while we only had seven raters, each rater could apply multiple codes to the same response, and many responses included both positive and negative themes; as a result, the denominator in this calculation is often .) Responses to question 36 (“How do you feel about reading scientific research papers?”) often contained a mix of positive and negative codes, leading to a broader range of written scores. Questions 41 (capable) and 42 (belonging) are more straightforward (i.e., only coded with ”yes,” ”no,” and ”maybe” for this analysis), and as such only have written scores of , , or .
Written response scores tend to track Likert scores for each of the questions, as computed with the Spearman correlation coefficient (where a score of 0 indicates no correlation, and a monotonic correlation). However, the correlation between the two scores is smaller in the post-survey than in the pre-survey for all three questions. (In the pre-survey, the coefficients are , , and for questions 39, 41, and 42, respectively; in the post-survey, these are , , and .) We also found that many students’ quantitative scores changed more substantially than their qualitative ones; typically, their Likert scale responses were more positive while their written response scores did not change or showed only minor change. Additionally, students occasionally responded to Likert score questions positively (i.e., ) while using neutral or negative language in the corresponding written response question. This inconsistency reflects the greater nuance available to students in written response.
VI Discussion
Both students and educators expressed positive feelings towards the accessible summary reading assignment in their direct feedback, and our data show positive change in all major categories probed by the quantitative survey questions (see Table 1), as summarized in Section III. The strongest effect sizes were measured for ability with jargon and main takeaways.
@@ VI Discussion
The amount of students whose scores began and remained high (as shown in Figure 4 for the quantitative data, and Figure 3 for one of the qualitative questions) highlights a limitation in our analysis: students whose initial responses lay to one extreme or another (the “floor and ceiling effects” [60]), i.e., students whose initial averages in a category were close to 1 or 7. A student with an initial average of 2 could not exhibit a decrease of more than after experiencing the interventions, while a student with an initial average of 6 could not exhibit an increase of more than . Additionally, decreases or lack of change in confidence – as in Q38 – may also be attributed to the fact that students’ “measuring stick” for themselves has changed since the beginning of the course, as they have gained knowledge and practice [61, 62]. It is also important to note that the timing of post-survey administration varied between instructors in the study (some administered it immediately after completion of the activity, some provided a window of time such as a week after the activity to complete the survey on their own time), and this may have had an effect on student responses, providing them more time in which to gain said knowledge and practice. Furthermore, we only completed the treatment once, and it may take repeated exposures (i.e. more practice with the assignments) for benefits to be realized.
@@
Qualitative responses from students showed a more mixed and nuanced view of their development from before to after experiencing the intervention. In response to questions about feelings while reading research, students less often mentioned feelings of intimidation and unpreparedness after the lesson, while mentioning more often that papers feel important/worthwhile and they are confident in their abilities.
@@
For a majority of students, the main takeaway from the lesson was either a greater understanding of the field of astronomy – e.g. how research works, what scientists do, how research publications come to be – or an increased awareness of existing resources and the need for additional resources to support accessibility in the field. Nearly all students’ takeaways were positive and in line with the learning objectives from the assignment template, also including increased content knowledge and increased confidence with course material. Although % of students claim their feelings about astronomy have not changed as a result of the activity, those who do claim to have changed are largely positive.
@@
Students were generally consistent in their qualitative and quantitative responses for the pre-survey, but less so in the post-survey. We argue that the greater nuance available in written responses meant that students were more likely to express a mix of positive and negative feelings – with our qualitative scoring model for consistency analysis, this could translate to more neutral or negative scores, even if their overall feelings (as reflected through a collection of Likert scale questions) showed positive change. Additionally, students and instructors were largely consistent in their feedback; instructors reported feeling that their students had gained understanding and perspective from the activities, and many students reported either improved or reinforcedconfidence in their abilities.
VII Conclusions & Recommendations